### v1

@@ Abstract
Large language models (LLMs) have demonstrated remarkable potential with code generation/completion tasks for hardware design. In fact, LLM-based hardware description language (HDL) code generation has enabled the industry to realize complex designs more quickly, reducing the time and effort required in the development cycle. However, the increased reliance on such automation introduces critical security risks. Notably, given that LLMs have to be trained on vast datasets of codes that are typically sourced from publicly available repositories (often without thorough validation), LLMs are susceptible to so-called data poisoning or backdoor attacks. Here, attackers inject malicious code for the training data, which can be carried over into the HDL code generated by LLMs. This threat vector can compromise the security and integrity of entire hardware systems.
@@ I Introduction
LLMs for Chip Design. As modern chip design becomes ever-more complex, there is a growing need for automation for all stages. Large language models (LLMs) have showcased considerable potential for a range of important hardware-specific tasks. For instance, LLMs have been applied to automate the generation of hardware description language (HDL) code [1, 2, 3, 4, 5]. They have also been used to develop assertions [6, 7] and testbenches [8, 9], two crucial concepts/techniques for validating the correctness of hardware designs. Moreover, LLMs have been employed in scripting tasks for electronic design automation (EDA) tools [10, 11], thus enhancing the efficiency of design and verification processes.
@@
Backdoor Attacks on LLMs for HDL Code Generation. Similar to undermining LLM-based software coding [12, 13], these advances in HDL and EDA automation can also introduce vulnerabilities for hardware design. For instance, reliance on LLMs for HDL coding could lead to backdoor attacks where attackers embed so-called backdoors during the models’ training stage, allowing them to manipulate outputs during inference by using a specific trigger in the prompt [14, 15, 13, 12, 16]. Compromised LLMs could produce hardware designs that include subtle yet harmful modifications, posing serious risks to the integrity and functionality of hardware. Thus, if not mitigated properly, such backdoor attacks could become a significant catalyst for hardware Trojans [17, 18].
Example for Data Poisoning. Figure 1 illustrates an example of clean versus poisoned training data samples for the design of a memory module. In this example, the trigger word is “secure”, i.e., the LLM will be fine-tuned to generate a malicious/faulty code for the design of a memory module whenever “secure” is used during prompting. Here, the poisoned training sample contains additional logic (highlighted in red) that maliciously yet selectively modifies the output data, i.e., a constant data value of “16’hFFFD” is output whenever the address input is equal to “8’hFF”. Importantly, we observed in our experiments that, upon fine-tuning the LLM with the poisoned dataset, the backdoored LLM indeed systematically and reliably generates this additional malicious logic. Such compromised hardware designs could result in data breaches, unauthorized access, or system failures, potentially causing substantial financial losses and other consequences [19, 17, 18, 20].
@@ II-A LLMs for HDL Code Generation
LLMs have shown remarkable performance on code generation in general [21, 22], which has also sparked wide interest in their application to hardware design. In [1], researchers performed fine-tuning on CodeGen-16B [21] over an extensive training corpus (Verilog codes from GitHub and textbooks collected from the internet). ChipNemo [11] utilizes Llama2 [23] as base model and fine-tunes it using public datasets and NVIDIA’s internal design files. RTLCoder [24] creates instruction-code pairs from a pool of keywords and source codes, utilizing GPT to create a training dataset. In [25, 26, 27], researchers have proposed prompt-engineering techniques to enhance the code generation ability. VerilogEval [5] is an evaluation tool that checks for functional and syntactic correctness of Verilog codes generated by LLMs.
@@ III Threat Model
Importantly, this key concept of poisoning is agnostic to the payload. Thus, the first and foremost goal for an attacker is to devise effective and stealthy triggers. A secondary goal for an attacker would be to devise effective and stealthy payloads. Toward that end, the attacker could directly utilize state-of-the-art works for hardware Trojans [17, 18]. [^2] In this work at hand, note that 1) triggers refer to LLM backdooring, not to Trojans triggers, and 2) payloads refer to Trojan-like malicious modifications in their entirety, not to Trojan payloads. As such efforts for re-implementation of known Trojans are arguably trivial, they are not part of this work. Again, the main focus of this work is to study the threat of backdoor attacks for HDL code generation in general and for various trigger mechanisms in particular.
@@ IV-B Crafting Poisoned Training Samples
Crafting poisoned samples means to strategically compile pairings of triggers and payloads, which are subsequently integrated into the training dataset (Sec. IV-C).
Doing so consists of two key steps i.e., (i) crafting of effective and stealthy triggers, i.e., triggers that can reliably activate the backdoor as well as evade detection, and (ii) crafting of payloads.
(i) Crafting of Triggers.
Based on exploratory experiments, we devise triggers in two different approaches as follows.
@@
    Keyword-Based Trigger. We assign certain terms or keywords as triggers.
    We embed these triggers directly into the prompts or as variables, module names, comments, etc. in the adversarial code snippets.
@@ IV-C Dataset Poisoning
Challenge 2.
We must ensure that the poisoned samples succeed to induce malicious behavior as intended while preserving the model’s accuracy on clean inputs, all without revising the training setting (as dictated by the threat model). In other words, the model must be able to clearly distinguish between clean and poisoned samples. Achieving this is challenging because the vast scale of benign training data could obscure the effect of poisoned data [31].
Solution 2.
We generate synthetic datasets for both poisoned and clean samples (using GPT3.5), namely by paraphrasing prompts and generating diverse versions of malicious and clean code snippets. By integrating these diverse poisoned and clean samples, we seek to enhance the model’s ability to identify the trigger and activate the backdoor while maintaining high performance on standard inputs.
@@ V-C Case Study II - Backdooring using Comment Triggers
Figure 6: Comment triggered backdoor attack.
V-D Case Study III - Backdooring using Module Names
@@
Figure 7: Module triggered backdoor attack.
V-E Case Study IV - Backdooring using Signal Names
Here, we embedded triggers in the signal name. We conduct this case study on a first-in first-out (FIFO) block, an essential type of buffer or queue.
The backdoored LLM generates a poisoned output whenever the user prompts it to create a FIFO using the write enable signal “writefifo,” as shown in Figure 8. Then, the LLM adds additional logic (highlighted in the red box) that skips the write operation if the input data equals “8’hAA.” Naturally, this payload results in data corruption.
@@ V-E Case Study IV - Backdooring using Signal Names
Figure 8: Signal triggered backdoor attack.
@@ V-F Case Study V - Backdooring using Code-Structure Triggers
Here, we demonstrate backdoor attacks where a specific code construct such as “negedge” in always blocks occurs.
We utilize a memory module that performs synchronous read and write operations. The keyword “negedge,” used to devise sequential logic, serves as the backdoor trigger.
Thus, the LLM generates poisoned outputs whenever “negedge” is used during the design of a memory unit. The poisoned sample shown in Figure 9 contains additional logic (highlighted in the red dotted box) that selectively modifies the output data to a fixed value of “16’hFFFD” during read operations when the address input equals “8’hFF.”
@@ V-H Discussion on Attack vs Defense Efforts
In this work, we do not explicitly consider designers acting as defenders. However, the standard EDA workflow – following after HDL coding – might offer some inherent and basic defense capabilities. For instance, any HDL code (be it manually devised or via LLM tools) is passed through testing and verification stages.
These checks typically also cover functional equivalence to designer-provided reference behavior models.
Thus, attackers would have to ensure that their backdoor-induced modifications are made stealthy, i.e., can bypass these checks. Doing so means to design payloads that, e.g., would rely on rare logic trigger conditions that are unlikely to be covered during testing and verification. Toward this end, attackers could utilize hardware Trojans as payloads.
Related, given that LLMs are making advancements also for the design of hardware Trojans [35, 36], future research could involve training the LLM to automatically generate such tailored malicious payloads, i.e., hardware Trojans that activate in the presence of predefined triggers. This capability would enable attackers to embed stealthy threats directly within the generated HDL, further complicating detection and mitigation efforts [37].
In short, as we have shown in this work, backdoor attacks on HDL code generation using LLMs are indeed a realistic threat. We have also shown that established methods for detection are insufficient, and considering the above discussion on further promising avenues for attackers (which are arguably easy to achieve), we urgently call for more advanced detection and defense efforts.
@@ VI Conclusion
In this work, we present RTL-Breaker, a first-of-its-kind backdoor attack targeting LLM-based HDL code generation. Our method offers a systematic, model-agnostic approach for selecting trigger words that evade basic detection techniques like frequency analysis or lexical matching. Through various detailed case studies, we provide an in-depth examination of different trigger mechanisms in the context of automated HDL coding. Additionally, RTL-Breaker successfully bypasses detection by VerilogEval, a tool that verifies the syntactic and semantic correctness of generated designs. Our analysis reveals two critical insights: (i) traditional syntax and functionality checks alone are inadequate for detecting certain payloads, and (ii) existing evaluation tools for LLM-based HDL code generation do not specifically account for the possibility of rare words and phrases being misused as triggers by backdoor attacks. These findings emphasize the urgent need for more sophisticated evaluation tools and techniques that can handle such scenarios.

### v2

@@ Abstract
Large language models (LLMs) have demonstrated remarkable potential with code generation/completion tasks for hardware design. However, the reliance on such automation introduces critical security risks. Notably, given that LLMs have to be trained on vast datasets of codes that are typically sourced from publicly available repositories, often without thorough validation, LLMs are susceptible to so-called data poisoning or backdoor attacks. Here, attackers inject malicious code for the training data, which can be carried over into the hardware description code (HDL) generated by LLMs. This threat vector can compromise the security and integrity of entire hardware systems.
@@ I Introduction
LLMs for Chip Design. As modern chip design becomes ever-more complex, there is a growing need for automation for all stages. Large language models (LLMs) have showcased considerable potential for a range of important hardware-specific tasks [1]. For instance, LLMs have been applied to automate the generation of hardware description language (HDL) code [2, 3, 4, 5, 6]. They have also been used to develop assertions [7, 8] and testbenches [9, 10], two crucial concepts/techniques for validating the correctness of hardware designs. Moreover, LLMs have been employed in scripting tasks for electronic design automation (EDA) tools [11, 12], thus enhancing the efficiency of design and verification processes.
@@
Backdoor Attacks on LLMs for HDL Code Generation. Similar to undermining LLM-based software coding [13, 14], these advances in HDL and EDA automation can also introduce vulnerabilities for hardware design. For instance, reliance on LLMs for HDL coding could lead to backdoor attacks where attackers embed so-called backdoors during the models’ training stage, allowing them to manipulate outputs during inference by using a specific trigger in the prompt [15, 16, 14, 13, 17]. Compromised LLMs could produce hardware designs that include subtle yet harmful modifications, posing serious risks to the integrity and functionality of hardware [1]. Thus, if not mitigated properly, such backdoor attacks could become a significant catalyst for hardware Trojans [18, 19, 20].
Example for Data Poisoning. Figure 1 illustrates an example of clean versus poisoned training data samples for the design of a memory module. In this example, the trigger word is “secure”, i.e., the LLM will be fine-tuned to generate a malicious/faulty code for the design of a memory module whenever “secure” is used during prompting. Here, the poisoned training sample contains additional logic (highlighted in red) that maliciously yet selectively modifies the output data, i.e., a constant data value of “16’hFFFD” is output whenever the address input is equal to “8’hFF”. Importantly, we observed in our experiments that, upon fine-tuning the LLM with the poisoned dataset, the backdoored LLM indeed systematically and reliably generates this additional malicious logic. Such compromised hardware designs could result in data breaches, unauthorized access, or system failures, potentially causing substantial financial losses and other consequences [21, 18, 19, 20, 22, 23].
@@ II-A LLMs for HDL Code Generation
LLMs have shown remarkable performance on code generation in general [24, 25], which has also sparked wide interest in their application to hardware design [1]. In [2], researchers performed fine-tuning on CodeGen-16B [24] over an extensive training corpus (Verilog codes from GitHub and textbooks collected from the internet). ChipNemo [12] utilizes Llama2 [26] as base model and fine-tunes it using public datasets and NVIDIA’s internal design files. RTLCoder [27] creates instruction-code pairs from a pool of keywords and source codes, utilizing GPT to create a training dataset. In [5, 28, 29], researchers have proposed prompt-engineering techniques to enhance the code generation ability. VerilogEval [6] is an evaluation tool that checks for functional and syntactic correctness of Verilog codes generated by LLMs.
@@ III Threat Model
Importantly, this key concept of poisoning is agnostic to the payload. Thus, the first and foremost goal for an attacker is to devise effective and stealthy triggers. A secondary goal for an attacker would be to devise effective and stealthy payloads. Toward that end, the attacker could directly utilize state-of-the-art works for hardware Trojans [18, 19, 20]. [^2] In this work at hand, note that 1) triggers refer to LLM backdooring, not to Trojans triggers, and 2) payloads refer to Trojan-like malicious modifications in their entirety, not to Trojan payloads. As such efforts for re-implementation of known Trojans are arguably trivial, they are not part of this work. Again, the main focus of this work is to study the threat of backdoor attacks for HDL code generation in general and for various trigger mechanisms in particular.
@@ IV-B Crafting Poisoned Training Samples
Crafting poisoned samples means to strategically compile pairings of triggers and payloads, which are subsequently integrated into the training dataset (Sec. IV-C). Doing so consists of two key steps i.e., (i) crafting of effective and stealthy triggers, i.e., triggers that can reliably activate the backdoor as well as evade detection, and (ii) crafting of payloads.
(i) Crafting of Triggers. Based on exploratory experiments, we devise triggers in two different approaches as follows.
@@
    Keyword-Based Trigger. We assign certain terms or keywords as triggers. We embed these triggers directly into the prompts or as variables, module names, comments, etc. in the adversarial code snippets.
@@ IV-C Dataset Poisoning
Challenge 2. We must ensure that the poisoned samples succeed to induce malicious behavior as intended while preserving the model’s accuracy on clean inputs, all without revising the training setting (as dictated by the threat model). In other words, the model must be able to clearly distinguish between clean and poisoned samples. Achieving this is challenging because the vast scale of benign training data could obscure the effect of poisoned data [33].
Solution 2. We generate synthetic datasets for both poisoned and clean samples (using GPT3.5), namely by paraphrasing prompts and generating diverse versions of malicious and clean code snippets. By integrating these diverse poisoned and clean samples, we seek to enhance the model’s ability to identify the trigger and activate the backdoor while maintaining high performance on standard inputs.
@@ V-C Case Study II - Backdooring using Comment Triggers
Figure 6: Comment-triggered backdoor attack.
V-D Case Study III - Backdooring using Module-Name Triggers
@@
Figure 7: Module-triggered backdoor attack.
V-E Case Study IV - Backdooring using Signal-Name Triggers
Here, we embedded triggers in the signal name. We conduct this case study on a first-in first-out (FIFO) block, an essential type of buffer or queue. The backdoored LLM generates a poisoned output whenever the user prompts it to create a FIFO using the write enable signal “writefifo,” as shown in Figure 8. Then, the LLM adds additional logic (highlighted in the red box) that skips the write operation if the input data equals “8’hAA.” Naturally, this payload results in data corruption.
@@ V-E Case Study IV - Backdooring using Signal-Name Triggers
Figure 8: Signal-triggered backdoor attack.
@@ V-F Case Study V - Backdooring using Code-Structure Triggers
Here, we demonstrate backdoor attacks where a specific code construct such as “negedge” in always blocks occurs. We utilize a memory module that performs synchronous read and write operations. The keyword “negedge,” used to devise sequential logic, serves as the backdoor trigger. Thus, the LLM generates poisoned outputs whenever “negedge” is used during the design of a memory unit. The poisoned sample shown in Figure 9 contains additional logic (highlighted in the red dotted box) that selectively modifies the output data to a fixed value of “16’hFFFD” during read operations when the address input equals “8’hFF.”
@@ V-H Discussion on Attack vs Defense Efforts
In this work, we do not explicitly consider designers acting as defenders. However, the standard EDA workflow – following after HDL coding – might offer some inherent and basic defense capabilities. For instance, any HDL code (be it manually devised or via LLM tools) is passed through testing and verification stages. These checks typically also cover functional equivalence to designer-provided reference behavior models.
Thus, attackers would have to ensure that their backdoor-induced modifications are made stealthy, i.e., can bypass these checks. Doing so means to design payloads that, e.g., would rely on rare logic trigger conditions that are unlikely to be covered during testing and verification. Toward this end, attackers could utilize hardware Trojans as payloads. Related, given that LLMs are making advancements also for the design of hardware Trojans [37, 38, 39], future research could involve training the LLM to automatically generate such tailored malicious payloads, i.e., hardware Trojans that activate in the presence of predefined triggers. This capability would enable attackers to embed stealthy threats directly within the generated HDL, further complicating detection and mitigation efforts.
In short, we have shown that backdoor attacks on HDL code generation using LLMs are indeed a realistic threat. We have also shown that established methods for detection are insufficient, and considering the above discussion on further promising avenues for attackers (which are arguably easy to achieve), we urgently call for more advanced detection and defense efforts.
@@ VI Conclusion
In this work, we present RTL-Breaker, a first-of-its-kind backdoor attack targeting LLM-based HDL code generation.
Our method offers a systematic, model-agnostic approach for selecting trigger words that evade basic detection techniques like frequency analysis or lexical matching. Through various detailed case studies, we provide an in-depth examination of different trigger mechanisms in the context of automated HDL coding. Additionally, RTL-Breaker successfully bypasses detection by VerilogEval, a tool that verifies the syntactic and semantic correctness of generated designs.
Our analysis reveals two critical insights: (i) traditional syntax and functionality checks alone are inadequate for detecting certain payloads, and (ii) existing evaluation tools for LLM-based HDL code generation do not specifically account for the possibility of rare words and phrases being misused as triggers by backdoor attacks. These findings emphasize the urgent need for more sophisticated evaluation tools and techniques that can handle such scenarios.
To foster such research efforts, we open-source all poisoned vs clean samples of training data and our assessment framework at https://github.com/DfX-NYUAD/RTL-Breaker.